A company produces and sells two products. Data concerning those products for the most recent month appear below:

Product Q91I Product J53Z

sales $15,000 $11,000

Variable expenses $5,850 $5,070

Fixed expenses for the entire company were $13,980.

a. Determine the overall break-even for the company. Show your Work.

b. If the sales mix shifts toward Product Q91I with no change in total sales, what will happen to the break-even point for the company? Explain.
Answer #1

WN- calculation of entire PV ratio-

total sales = 15000+11000 = 26000

total variable expense = 5850+5070 = 10920

Profit = 26000-10920 = 15080

PV ratio = 15080/26000*100 = 58%

a. overall break even = fixed expense/PV ratio

=13980/58%

= 24103.45

b. total sales = 26000

variable expense ratio od product Q91I = 5850/15000*100 = 39%

PV ratio of productvQ91I = 61%

break even = 13980/61% = 22918.03